Chris reached into his pocket and pulled out twenty coins. the coins consisted of nickels, dimes, and quarters. If the number of dimes is two less than the number of nickels, and if the total value of the coins is $2.50, then how many of each coin did he have in his pocket?
Answer by TimothyLamb(4379)

---
x + y + z = 20
5x + 10y + 25z = 250
y = x - 2
---
put the system of linear equations into standard form
---
x + y + z = 20
5x + 10y + 25z = 250
x - y = 2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= nickels = 8
y= dimes = 6
z= quarters = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php